# Adversarial Fine-tuning in Offline-to-Online Reinforcement Learning for Robust Robot Control

### v1

@@ 1 Introduction
Our key insight is that conservative offline RL and action-space robustness are fundamentally incompatible. Conservative methods constrain policies to dataset actions to prevent extrapolation errors, yet robustness to action perturbations requires learning from the out-of-distribution samples that these constraints prohibit. This incompatibility is illustrated in Figure 1. In the left panel, a conservatively trained policy produces an action distribution (blue) that aligns with the offline dataset (dashed) under normal conditions. However, when action perturbations occur, the executed actions shift into out-of-distribution regions (red) where the policy fails. The right panel illustrates that if the policy could learn to produce compensatory actions that keep the executed action distribution consistent across both normal and perturbed conditions, robustness would be achieved (blue and red overlap).
@@
However, existing work on robustness in offline RL has focused primarily on distributional shifts in the state-space, leaving this fundamental incompatibility largely unaddressed. Most prior work addresses state perturbations (e.g., sensor noise) through smoothness penalties and output stabilization [14, 15, 16], which remain compatible with conservative constraints. In contrast, action perturbations (e.g., actuator faults) have been studied far less. These disturbances directly modify executed actions, and recent work shows that offline RL is particularly vulnerable to such perturbations [17], as datasets typically lack perturbation-aware transitions while conservative constraints prevent necessary exploration.
@@ 2.2 Robust RL
In reinforcement learning for robot control, robustness is crucial for handling uncertainties such as sensor noise, matched physical parameter, unexpected terrain variations, or mechanical failures [21, 22, 23, 24, 25]. A common approach is to introduce perturbations during training to improve policy generalization across conditions. Domain randomization [26, 27, 28] improves adaptability by randomizing environment parameters such as friction, while adversarial training [23, 29, 30, 31, 32] improves robustness by exposing policies to adversarial perturbations.
@@ 2.3 Offline-to-Online RL
A major challenge in offline-to-online RL is performance degradation in the early phase of fine-tuning, caused by distributional shifts between offline data and newly collected online experiences [10, 11]. Various approaches have been proposed to mitigate this issue. Representative examples include adaptive or relaxed policy constraints [10, 35, 12], data selection methods that prioritize near-on-policy samples [11], dynamic switching among multiple policies depending on the state [13, 34], and exploration strategies combined with meta-adaptation [33].
This study leverages offline-to-online RL to improve robustness against action-space perturbations. Such perturbations represent discrepancies between training and deployment environments. To address this challenge, the policy is first pretrained offline and then fine-tuned in an online environment where perturbations are introduced. This approach enables policies to acquire robustness to action-space perturbations adaptively, which is difficult to achieve in purely offline settings.
@@ 3.2 Offline-to-Online RL
Some prior work [35, 12] retains the TD3+BC penalty during fine-tuning. However, when perturbations are introduced, the effectiveness of stored actions under perturbed dynamics is not guaranteed. In such cases, maintaining the constraint may hinder adaptation. Therefore, we discard the penalty during online fine-tuning to allow more flexible policy updates.
@@ 4 Proposed Method
This study addresses the problem of improving policy robustness against action-space perturbations in the offline RL setting. As discussed in Section 3.1, offline RL constrains policies to stay close to actions in the dataset, as shown in Equation (1). Thus, when the dataset lacks transitions with perturbed actions, it is fundamentally difficult to obtain a policy that adapts to perturbed environments. To solve this limitation, we adopt an offline-to-online approach, pretraining a policy on a clean offline dataset and fine-tuning it in online adversarial environments.
@@ 4.2 Adversarial Fine-tuning
$\boldsymbol{a}_{t}^{\prime}=\boldsymbol{a}_{t}+\boldsymbol{\delta}_{a}\odot\boldsymbol{a}_{t}\quad\text{with probability }q,$ (8)
@@ 4.3.1 Linear curriculum
where $c$ is a constant step size. Although straightforward, overly large $c$ may lead to excessive exposure to perturbations too early, which can degrade nominal performance under normal (unperturbed) conditions (see Appendix A).
@@ 4.3.2 Adaptive curriculum
The exponential moving average $\bar{R}_{n}$ provides a stable estimate of performance trends by filtering out short-term fluctuations. The parameters $\eta$ and $\beta$ control the adaptation dynamics: a smaller $\beta$ captures long-term stability, while a larger $\beta$ responds more quickly to recent changes. A larger $\eta$ leads to stronger adaptation, whereas a smaller $\eta$ results in smoother updates. By adjusting these parameters, the adaptive curriculum gradually modifies $q$ to improve robustness without causing instability. The details of these parameter settings are provided in Appendix A.
@@ 5.1 Experimental Setup
Fine-Tuning Settings Fine-tuning is conducted for 1 million steps using the TD3 algorithm. Each configuration is trained with 5 independent runs. The perturbation probability $q$ and the offline data ratio $r_{\text{off}}$ are set to $q=0.5$, $r_{\text{off}}=0$ in the Hopper environment. In the HalfCheetah and Ant environments, $q$ is set to $0.1$ and $r_{\text{off}}$ to $0.1$.
@@
In the normal condition, no perturbation is applied ($\boldsymbol{\delta}=\boldsymbol{0}$). For the random perturbation condition, $\boldsymbol{\delta}_{r}$ is sampled from a uniform distribution over $\epsilon,\epsilon]^{N_{a}}$, where $\epsilon$ controls the perturbation strength and $N_{a}$ denotes the action dimension. In the Hopper-v2 and HalfCheetah-v2 environments, the magnitudes of $\boldsymbol{\delta}_{r}$ and $\boldsymbol{\delta}_{a}$ are fixed at $\epsilon=0.3$ during both training and evaluation. In the Ant-v2 environment, the perturbation magnitude is fixed at $\epsilon=0.5$.
@@
Robustness Evaluation We evaluate the robustness of fine-tuned policies on a forward walking task in a legged robot environment. Each policy is evaluated under three distinct conditions: normal (i.e., no perturbation), with random perturbations, and with adversarial perturbations. In the random and the adversarial perturbation condition, the perturbation probability is set to $q=1$. The evaluation metric is the D4RL-normalized episodic reward over 100 episodes in each condition. All experiment results are averaged across the 5 independent runs.
@@ 5.2 Robustness Evaluation of Fine-tuned Models against Baselines
Table 1: Fine-tuning improves robustness and achieves the highest scores, in contrast to conservative offline methods that drop to sub-zero rewards under adversarial perturbations. Values are reported as D4RL-normalized episodic reward ($\text{mean}\pm\text{std}$). Offline results are from [17]. Fully Online denotes policies trained from scratch under adversarial perturbations. Fine-tuned denotes pretrained offline policies further trained under the specified perturbation condition. Bold numbers indicate the best performance for each evaluation condition.
Environments
Evaluation
Conditions
Offline
Fully Online
Fine-tuned
@@
Hopper-v2 Normal $\bf{111.1\pm 2.9}$ $0.9\pm 39.4 88$$8.5\pm 40.3$ $102.5\pm 20.1$ $84.3\pm 30.0$
Random $62.9\pm 42.8$ $77.2\pm 40.2$ $78.5\pm 41.1$ $\bf{95.8\pm 24.4}$ $89.3\pm 24.6$
Adversarial $13.7\pm 2.4$ $57.0\pm 33.1$ $16.7\pm 1.6$ $57.4\pm 27.1$ $\bf{83.5\pm 17.9}$
HalfCheetah-v2 Normal $\bf{96.6\pm 3.4}$ $34.2\pm 10.8$ $92.7\pm 4.2$ $93.7\pm 6.2$ $89.7\pm 7.4$
Random $75.0\pm 18.8$ $30.5\pm 10.2$ $73.4\pm 14.4$ $\bf{82.7\pm 10.8}$ $73.8\pm 13.7$
@@
Ant-v2 Normal $119.9\pm 39.8$ $34.4\pm 12.3$ $127.7\pm 31.5$ $\bf{134.2\pm 25.4}$ $130.0\pm 18.5$
Random $86.6\pm 43.3$ $31.8\pm 12.8$ $103.4\pm 36.3$ $\bf{104.6\pm 33.8}$ $99.6\pm 32.8$
Adversarial $21.0\pm 25.9$ $24.0\pm 16.2$ $32.9\pm 32.2$ $56.2\pm 25.4$ $\bf{91.6\pm 11.9}$
@@
Table 1 compares adversarial fine-tuning against two baselines: offline-only training and fully online training from scratch. Offline-only policies fail under action perturbations due to their lack of exposure to distribution shifts during training. For example, in Ant-v2 under adversarial conditions, adversarial fine-tuning achieves a score of 91.6, compared to -21.0 for offline training and 24.0 for fully online training. A similar trend is observed in Hopper-v2, where the offline score drops to 13.7 while the fine-tuned score reaches 83.5. These results highlight that adversarially fine-tuned policies are far more robust to action-space perturbations than policies trained solely offline. The results also show that matching the fine-tuning condition to the evaluation condition (e.g., Fine-tuned (Adversarial) for adversarial evaluation) yields the best performance.
@@
These results demonstrate that adversarial fine-tuning effectively combines the robustness advantages of online training with the sample efficiency of offline pretraining. This approach provides a practical solution for deploying RL policies in real-world robotic systems where actuator perturbations are inevitable.
@@ 5.3 Improving Adversarial Robustness via Curriculum Fine-tuning
Table 2 compares the three strategies across training durations and evaluation conditions. At 1M steps under adversarial conditions, $q_{\text{lin}}$ shows inconsistent performance: lower than $q_{\text{fix}}$ in Hopper-v2 (30.7 vs 83.5) and Ant-v2 (84.6 vs 91.6), but comparable in HalfCheetah-v2 (58.7 vs 55.3). In contrast, $q_{\text{ada}}$ consistently outperforms $q_{\text{fix}}$ across all three environments (Hopper-v2: 87.9 vs 83.5; HalfCheetah-v2: 63.4 vs 55.3; Ant-v2: 98.2 vs 91.6), demonstrating more reliable early-stage robustness. By 3M steps, the two curricula exhibit distinct trade-offs. $q_{\text{lin}}$ improves adversarial performance in some environments (Ant-v2: 84.6 at 1M to 117.7 at 3M) but degrades normal performance (Hopper-v2: 95.1 to 76.5; HalfCheetah-v2: 91.7 to 88.7; Ant-v2: 125.0 to 123.7). Figure 5 confirms this overfitting phenomenon, showing $q$ increasing continuously throughout training. In contrast, $q_{\text{ada}}$ maintains or improves normal performance across environments (Hopper-v2: 100.2 to 97.5; HalfCheetah-v2: 89.0 to 93.0; Ant-v2: 124.8 to 130.8) while achieving comparable adversarial robustness (Ant-v2: 118.5 vs 117.7). This occurs because the adaptive schedule reduces perturbation pressure when performance drops, preventing unnecessary difficulty escalation.
This demonstrates that performance-based adaptation successfully avoids the overfitting-robustness trade-off inherent in fixed scheduling. Moreover, it suggests that dynamic curriculum mechanisms can offer a principled way to balance robustness and stability in broader reinforcement learning settings.
@@
Environments
Evaluation
Conditions
$q_{\text{fix}}$ $q_{\text{lin}}$ $q_{\text{ada}}$
@@
Figure 5: The trajectories of the perturbation probability $q$ illustrate the key behavioral difference between the two curricula. The adaptive strategy ($q_{\text{ada}}$, red) adjusts the growth of $q$ in response to policy performance, preventing the relentless increase seen in the linear strategy ($q_{\text{lin}}$, blue).

### v2

@@ 1 Introduction
Our key insight is that conservative offline RL and action-space robustness are incompatible. Conservative methods constrain policies to dataset actions to prevent extrapolation errors, yet robustness to action perturbations requires learning from the out-of-distribution samples that these constraints prohibit. This incompatibility is illustrated in Figure 1. In the left panel, a conservatively trained policy produces an action distribution (blue) that aligns with the offline dataset (dashed) under normal conditions. However, when action perturbations occur, the executed actions shift into out-of-distribution regions (red) where the policy fails. The right panel illustrates that if the policy could learn to produce compensatory actions that keep the executed action distribution consistent across both normal and perturbed conditions, robustness would be achieved (blue and red overlap).
@@
However, existing work on robustness in offline RL has focused primarily on distributional shifts in the state-space, leaving this incompatibility largely unaddressed. Most prior work addresses state perturbations (e.g., sensor noise) through smoothness penalties and output stabilization [39, 25, 28], which remain compatible with conservative constraints. In contrast, action perturbations (e.g., actuator faults) have been studied far less. These disturbances directly modify executed actions, and recent work shows that offline RL is particularly vulnerable to such perturbations [2], as datasets typically lack perturbation-aware transitions while conservative constraints prevent necessary exploration.
@@ 2.2 Robust RL
Robustness under uncertainty has been widely studied not only in RL but also in other domains such as cybersecurity and remote sensing [38, 13], where models must operate reliably under environmental variability or adversarial conditions. Within RL, robustness has also been explored through ensemble-based approaches [6, 1, 17], which combine multiple models or representations to reduce sensitivity to input variability and distributional shifts. These approaches mainly aim to improve stability and generalization during training under uncertain inputs. Ensemble-based or representation-level approaches primarily address robustness to input variability during training. In contrast, the present study focuses on robustness against action-space perturbations that occur at execution time in control tasks.
In RL for robot control, robustness is crucial for handling uncertainties such as sensor noise, matched physical parameter, unexpected terrain variations, or mechanical failures [43, 33, 30, 29, 24]. A common approach is to introduce perturbations during training to improve policy generalization across conditions. Domain randomization [35, 5, 18] improves adaptability by randomizing environment parameters such as friction, while adversarial training [30, 34, 44, 32, 47] improves robustness by exposing policies to adversarial perturbations.
@@ 2.3 Offline-to-Online RL
A major challenge in offline-to-online RL is performance degradation in the early phase of fine-tuning, caused by distributional shifts between offline data and newly collected online experiences [27, 22]. Various approaches have been proposed to mitigate this issue. Representative examples include adaptive or relaxed policy constraints [27, 46, 45], data selection methods that prioritize near-on-policy samples [22], dynamic switching among multiple policies depending on the state [41, 37], and exploration strategies combined with meta-adaptation [14]. These offline-to-online RL methods primarily consider distributional shifts under nominal environments, where the mismatch arises from differences between offline data and newly collected online experiences during data collection. In contrast, this study focuses on action-space perturbations, representing discrepancies between training and deployment environments.
This study leverages offline-to-online RL to improve robustness against action-space perturbations. To address this challenge, the policy is first pretrained offline and then fine-tuned in an online environment where perturbations are introduced. This approach enables policies to acquire robustness to action-space perturbations adaptively, which is difficult to achieve in purely offline settings.
@@ 3.2 Offline-to-Online RL
Some prior work [46, 45] retains the TD3+BC penalty during fine-tuning. However, when perturbations are introduced, the effectiveness of stored actions under perturbed dynamics is not guaranteed. In such cases, maintaining the constraint may hinder adaptation. Therefore, we discard the penalty during online fine-tuning to allow more flexible policy updates. The Appendix C presents a comparison of final performance with and without the BC penalty.
3.3 Limitation of Conservative Offline RL under Action-space Perturbations
This section clarifies under what condition the conservative constraint in Eq. 1 can limit compensation for action-space perturbations introduced in Section 3.1.
Sufficient Condition
The following condition describes when the conservative constraint can prevent compensation.
Lemma 1 (Sufficient condition for constraint-induced limitation).
Suppose that maintaining performance under the perturbed environment requires intended actions that lie outside an $\epsilon$-neighborhood of $\pi_{\beta}(\cdot|\boldsymbol{s})$ for some states $\boldsymbol{s}$. Then any offline RL method enforcing $D(\pi(\cdot|\boldsymbol{s}),\pi_{\beta}(\cdot|\boldsymbol{s}))\leq\epsilon$ cannot generally be expected to learn such compensation from the dataset $\mathcal{D}$ alone.
Proof sketch.
The dataset $\mathcal{D}$ contains only clean transitions. It provides no supervision signal for perturbed outcomes. Meanwhile, the divergence constraint restricts policy updates within the $\epsilon$-neighborhood. If compensation requires a larger deviation, the update is explicitly restricted. Therefore, compensation cannot typically be learned under this constraint without additional online interaction. ∎
Interpretation and Practical Implication
This condition does not imply that robustness is impossible in all offline settings. If the offline dataset already covers the necessary compensatory actions, robustness may still be achievable offline. However, when perturbations are adversarial, they are designed to reduce performance under the nominal policy. Compensation then tends to require policy adjustments toward directions where the nominal policy performs poorly. Such directions often lie outside the high-density region of the behavior distribution. Even if they remain within the support, they may require shifting probability mass toward low-density regions.
such probability shifts are discouraged by conservative objectives. As a result, these conservative constraints are likely to become effectively binding along the directions required for compensation. This mechanism helps explain why additional online interaction can be necessary. Online data provide transitions under perturbed actions and can reduce the effective limitation imposed by purely conservative offline objectives.
@@ 4 Proposed Method
This study addresses the problem of improving policy robustness against action-space perturbations in the offline RL setting. As discussed in Section 3.1, offline RL constrains policies to stay close to actions in the dataset, as shown in Equation (1). Thus, when the dataset lacks transitions with perturbed actions, it is difficult to obtain a policy that adapts to perturbed environments. To solve this limitation, we adopt an offline-to-online approach, pretraining a policy on a clean offline dataset and fine-tuning it in online adversarial environments.
@@ 4.2 Adversarial Fine-tuning
$\tilde{\boldsymbol{a}}_{t}=\boldsymbol{a}_{t}+\boldsymbol{\delta}_{a}\odot\boldsymbol{a}_{t}\quad\text{with probability }q,$ (8)
@@ 4.3.1 Linear curriculum
where $c$ is a constant step size. Although straightforward, overly large $c$ may lead to excessive exposure to perturbations too early, which can degrade nominal performance under normal (unperturbed) conditions (see Appendix B).
@@ 4.3.2 Adaptive curriculum
The exponential moving average $\bar{R}_{n}$ provides a stable estimate of performance trends by filtering out short-term fluctuations. The parameters $\eta$ and $\beta$ control the adaptation dynamics: a smaller $\beta$ captures long-term stability, while a larger $\beta$ responds more quickly to recent changes. A larger $\eta$ leads to stronger adaptation, whereas a smaller $\eta$ results in smoother updates. By adjusting these parameters, the adaptive curriculum gradually modifies $q$ to improve robustness without causing instability. Theoretical analysis is provided in Appendix A, and the settings of $\eta$ and $\beta$ are detailed in Appendix B.
@@ 5.1 Experimental Setup
Fine-Tuning Settings Fine-tuning is conducted for 1 million steps using the TD3 algorithm. Each configuration is trained with 5 independent runs. The perturbation probability $q$ and the offline data ratio $r_{\text{off}}$ are set to $q=0.5$, $r_{\text{off}}=0$ in the Hopper environment. In the HalfCheetah and Ant environments, $q$ is set to $0.1$ and $r_{\text{off}}$ to $0.1$. Further evaluation under different values of $r_{\text{off}}$ is reported in the Appendix C.
@@
In the normal condition, no perturbation is applied ($\boldsymbol{\delta}=\boldsymbol{0}$). For the random perturbation condition, $\boldsymbol{\delta}_{r}$ is sampled from a uniform distribution over $\epsilon,\epsilon]^{N_{a}}$, where $\epsilon$ controls the perturbation strength and $N_{a}$ denotes the action dimension. In the Hopper-v2 and HalfCheetah-v2 environments, the magnitudes of $\boldsymbol{\delta}_{r}$ and $\boldsymbol{\delta}_{a}$ are fixed at $\epsilon=0.3$ during both training and evaluation. In the Ant-v2 environment, the perturbation magnitude is fixed at $\epsilon=0.5$. Adversarial perturbations used during evaluation are newly generated using the differential evolution algorithm and are independent of those used during fine-tuning.
@@
Robustness Evaluation We evaluate the robustness of fine-tuned policies on a forward walking task in a legged robot environment. Each policy is evaluated under three distinct conditions: normal (i.e., no perturbation), with random perturbations, and with adversarial perturbations. In the random and the adversarial perturbation condition, the perturbation probability is set to $q=1$. The evaluation metric is the D4RL-normalized episodic reward over 100 episodes in each condition. All experiment results are averaged across the 5 independent runs. Specifically, the D4RL-normalized episodic reward is calculated as
$\mathrm{Normalized\ Episodic\ Reward}=100\times\frac{\bar{R}_{\mathrm{episode}}-\mathrm{score}_{\mathrm{min}}^{\mathrm{D4RL}}}{\mathrm{score}_{\mathrm{max}}^{\mathrm{D4RL}}-\mathrm{score}_{\mathrm{min}}^{\mathrm{D4RL}}},$ (14)
where $\mathrm{score}_{\mathrm{min}}^{\mathrm{D4RL}}$ and $\mathrm{score}_{\mathrm{max}}^{\mathrm{D4RL}}$ are predefined for each environment in D4RL and correspond to the performance of a random policy and an expert policy, respectively.
@@ 5.2 Robustness Evaluation of Fine-tuned Models against Baselines
Table 1: Fine-tuning improves robustness and achieves the highest scores, in contrast to conservative offline methods that drop to sub-zero rewards under adversarial perturbations. Values are reported as D4RL-normalized episodic reward ($\text{mean}\pm\text{SE}$). Offline results are from [2]. Fully Online denotes policies trained from scratch under adversarial perturbations. Fine-tuned denotes pretrained offline policies further trained under the specified perturbation condition. Bold numbers indicate the best performance for each evaluation condition.
Environments Evaluation Conditions Offline Fully Online (Adversarial) Fine-tuned
@@
Hopper-v2 Normal $\bf{111.1\pm 0.1}$ $0.9\pm 1.8 88$$8.5\pm 1.8$ $102.5\pm 0.9$ $84.3\pm 1.3$
Random $62.9\pm 1.9$ $77.2\pm 1.8$ $78.5\pm 1.8$ $\bf{95.8\pm 1.1}$ $89.3\pm 1.1$
Adversarial $13.7\pm 0.1$ $57.0\pm 1.5$ $16.7\pm 0.1$ $57.4\pm 1.2$ $\bf{83.5\pm 0.8}$
HalfCheetah-v2 Normal $\bf{96.6\pm 0.2}$ $34.2\pm 0.5$ $92.7\pm 0.2$ $93.7\pm 0.3$ $89.7\pm 0.3$
Random $75.0\pm 0.8$ $30.5\pm 0.5$ $73.4\pm 0.6$ $\bf{82.7\pm 0.5}$ $73.8\pm 0.6$
@@
Ant-v2 Normal $119.9\pm 1.8$ $34.4\pm 0.6$ $127.7\pm 1.4$ $\bf{134.2\pm 1.1}$ $130.0\pm 0.8$
Random $86.6\pm 1.9$ $31.8\pm 0.6$ $103.4\pm 1.6$ $\bf{104.6\pm 1.5}$ $99.6\pm 1.5$
Adversarial $21.0\pm 1.2$ $24.0\pm 0.7$ $32.9\pm 1.4$ $56.2\pm 1.1$ $\bf{91.6\pm 0.5}$
@@
Based on the fine-tuning and perturbation settings defined in Section 5.1, we evaluate the proposed adversarial fine-tuning method by comparing it with several baseline and control settings. Specifically, we compare the proposed Fine-tuned (Adversarial) setting with: (1) Offline: Offline-only training without online interaction, (2) Fully Online (Adversarial): Training from scratch under adversarial perturbations, and (3) Fine-tuned (Normal) and Fine-tuned (Random): Non-adversarial fine-tuning settings. Fine-tuned (Normal) performs fine-tuning without applying perturbations, and Fine-tuned (Random) applies uniformly sampled random perturbations during fine-tuning. These comparisons allow us to isolate the contribution of offline pretraining and to distinguish adversarial perturbations from non-adversarial perturbation injection.
Table 1 presents the comparison of these settings across all evaluation conditions. Values in the table are reported as mean ± standard error (SE), where SE is computed over $500$ evaluation episodes across five independent runs. Offline-only policies fail under action perturbations due to their lack of exposure to distribution shifts during training. For example, in Ant-v2 under adversarial conditions, adversarial fine-tuning achieves a score of $91.6$, compared to 21.0-21.0 for offline training and $24.0$ for fully online training. A similar trend is observed in Hopper-v2, where the offline score drops to $13.7$ while the fine-tuned score reaches $83.5$. These results highlight that adversarially fine-tuned policies are far more robust to action-space perturbations than policies trained solely offline. Policies fine-tuned under the normal conditions exhibit severe performance degradation when evaluated under the adversarial perturbation condition (e.g., $16.7$ in Hopper-v2). This performance degradation indicates that robustness is not automatically acquired without explicit exposure to perturbed environments during training. Robustness is closely tied to the distribution encountered during fine-tuning. Therefore, performance cannot be maintained when the evaluation condition deviates from the training condition.
Policies fine-tuned under adversarial perturbations retain substantial performance under the random perturbation condition (e.g., $89.3$ in Hopper-v2). This retained performance suggests that learning under perturbed environments does not result in over-specialization to a specific perturbation pattern. Learning under perturbed environments also improves more fundamental control properties, such as action-output stability and compensatory control ability. These improvements contribute to a certain degree of general robustness. However, the best performance is achieved when the fine-tuning condition matches the evaluation condition. This dependency implies that the acquired robustness is not fully perturbation-independent. The acquired robustness includes perturbation-specific adaptation in addition to the shared stability improvement.
@@
In this study, sample efficiency is evaluated based on the number of online environment interactions required during fine-tuning.
From this perspective, these results demonstrate that adversarial fine-tuning effectively combines the robustness advantages of online training with the sample efficiency of offline pretraining. This approach provides a practical solution for deploying RL policies in real-world robotic systems where actuator perturbations are inevitable. To further examine the algorithmic generality of the proposed framework, we conduct additional experiments using Implicit Q-Learning (IQL) [19] pretraining; detailed results are provided in Appendix D. We further investigate how different perturbation generation mechanisms influence fine-tuning behavior; details are provided in Appendix E.
@@ 5.3 Improving Adversarial Robustness via Curriculum Fine-tuning
Table 2 compares the three strategies across training durations and evaluation conditions. At 11M steps under adversarial conditions, $q_{\text{lin}}$ shows inconsistent performance: lower than $q_{\text{fix}}$ in Hopper-v2 ($30.7$ vs $83.5$) and Ant-v2 ($84.6$ vs $91.6$), but comparable in HalfCheetah-v2 ($58.7$ vs $55.3$). In contrast, $q_{\text{ada}}$ consistently outperforms $q_{\text{fix}}$ across all three environments (Hopper-v2: $87.9$ vs $83.5$; HalfCheetah-v2: $63.4$ vs $55.3$; Ant-v2: $98.2$ vs $91.6$), demonstrating more reliable early-stage robustness. By 33M steps, the two curricula exhibit distinct trade-offs. Although $q_{\text{lin}}$ improves adversarial performance in some environments (Ant-v2: $84.6$ at 11M to $117.7$ at 33M), it degrades normal performance (Hopper-v2: $95.1$ to $76.5$; HalfCheetah-v2: $91.7$ to $88.7$; Ant-v2: $125.0$ to $123.7$). In contrast, $q_{\text{ada}}$ maintains or improves normal performance across environments (Hopper-v2: 100.2 to 97.5; HalfCheetah-v2: $89.0$ to $93.0$; Ant-v2: $124.8$ to $130.8$) while achieving comparable adversarial robustness (Ant-v2: $118.5$ vs $117.7$).
Figure 5 shows that under $q_{\text{lin}}$, the perturbation probability increases monotonically throughout training. In the linear schedule setting, the decline in the normal performance arises from a distributional imbalance induced by this monotonic increase in perturbation probability. As the perturbation probability increases, transitions stored in the replay buffer $\mathcal{R}$ become progressively dominated by data collected under the perturbed environment. Consequently, gradient updates are biased toward adaptation to noisy conditions, while stable control behaviors required under the normal environments receive relatively less reinforcement. This imbalance weakens nominal behaviors acquired during offline pretraining and leads to the observed performance degradation.
In contrast, the adaptive schedule adjusts $q$ according to policy performance. This adjustment prevents excessive concentration of transitions from the perturbed environment in the replay buffer and mitigates the distributional imbalance. This demonstrates that performance-based adaptation successfully avoids the stability-robustness trade-off inherent in fixed scheduling. Moreover, it suggests that dynamic curriculum mechanisms can offer a principled way to balance robustness and stability in broader reinforcement learning settings.
@@
Environments Evaluation Conditions $q_{\text{fix}}$ $q_{\text{lin}}$ $q_{\text{ada}}$
@@
Figure 5: The adaptive curriculum ($q_{\text{ada}}$), which adjusts the perturbation probability based on performance, mitigates the degradation in performance under the normal condition observed under the linear schedule ($q_{\text{lin}}$) and maintains robust performance across evaluation conditions. Results are reported for the fixed-$q$ setting ($q_{\text{fix}}$), linearly increasing strategy ($q_{\text{lin}}$), and adaptively adjusted strategy ($q_{\text{ada}}$) after $1\text{M}$, $2\text{M}$, and $3\text{M}$ training steps. Values indicate D4RL-normalized episodic rewards ($\text{mean}\pm\text{standard error}$) under normal, random, and adversarial perturbation conditions.
6 Limitations and Future Directions
This study has three main limitations. First, this study focuses on static multiplicative action perturbations within each episode. These perturbations represent continuous execution errors such as actuator degradation or gain mismatch. However, more complex forms of action uncertainty are not addressed. For example, action delays and actuator saturation change the timing between actions and state transitions or impose nonlinear constraints. Because these effects change the control dynamics, our framework cannot be directly applied without reformulation. Second, the proposed method addresses action-space (output-level) perturbations, while many prior robust RL approaches [43, 42] focus on state-space (input-level) perturbations. Since they affect different parts of the model, combining state-robust methods with the proposed action-space adaptation framework is a possible direction for future work. Finally, a more principled formulation that explicitly models the trade-off between nominal performance and adversarial robustness remains an important direction for future work. In the current framework, this balance is handled implicitly and may still depend on the specific task. An explicit multi-objective formulation [36, 3] could provide a clearer and more systematic understanding of this balance.